# Toward High-Performance LLM Serving: A Simulation-Based Approach for Identifying Optimal Parallelism

### v1

@@ 4.6. Beyond Identifying Optimal Execution Plan
While APEX is primarily developed to identify optimal parallel execution plans for LLM serving systems, its high simulation fidelity makes it applicable to other use cases as well. Below, we present two examples of applications.
4.6.1. APEX provides insights for building LLM serving platform
When a new hardware device is released, service providers can use APEX to estimate the expected performance improvements from adopting the new hardware by adjusting the hardware parameters of device cluster in APEX. For example, the performance boost of upgrading from Nvidia A100 to H100 GPUs can be projected by scaling the compute time and data transfer time according to their relative compute power and memory bandwidth. Using this approach, APEX estimates that upgrading from A100 to H100 results in a 1.79$\times$ improvement in TTFT and a 1.66$\times$ speedup in TPOT; this estimation closely aligns with actual results reported in (Patel et al., 2024), which shows a 1.85$\times$ improvement in TTFT and a 1.43$\times$ speedup in TPOT. APEX also allows users to freely scale hardware parameters (i.e., creating synthetic hardware) to estimate the hardware upgrade required to achieve a specific performance boost in LLM serving.
4.6.2. APEX provides insights for meeting SLOs
In addition to minimizing end-to-end serving latency, service providers must also meet various service-level objectives (SLOs), such as maintaining a time per output token (TPOT). APEX can assist in achieving these SLOs. A common strategy to meet latency requirements is to adjust the batch size by setting a maximum batch size constraint. APEX begins by simulating a subset of requests to determine an upper bound for the batch size, denoted as $m$. It then divides this upper bound into $n$ segments and simulates the request traces with various maximum batch size constraints, ranging from $\frac{1\times m}{n}$ to $\frac{n\times m}{n}$, where $n$ is determined heuristically. Figure 8 illustrates two examples using the Llama-3.1-70B and Mistral-Large models on the Creation dataset. For simplicity, only results for pure tensor parallelism are shown. However, similar evaluations are performed across all parallel execution plans, resulting in numerous design points to choose from. Assuming a service provider aims to decrease the TPOT to meet the SLO, they can use the results to estimate the necessary adjustment to the maximum batch size constraint. For instance, reducing the maximum batch size constraint from 16 to 8 results in an 18% TPOT improvement for the Llama model and a 14% improvement for the Mistral model. However, overly restricting the maximum batch size can negatively impact end-to-end latency, as illustrated in Figure 8, where the batch size constraint is set to 4.
@@ 4.6.2. APEX provides insights for meeting SLOs
5. Related Work
LLM Serving Systems: With the emergence of LLMs, numerous LLM serving systems have been proposed (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024; NVIDIA, 2024; Patel et al., 2024). Each system introduces innovations to address key challenges in LLM serving. Orca (Yu et al., 2022) introduced iteration-level batching, significantly improving serving throughput. This technique has since become a standard in LLM serving systems. vLLM (Kwon et al., 2023) proposed PagedAttention, an efficient method to manage KV cache memory usage, enabling more concurrent request batching. Sarathi-Serve (Agrawal et al., 2024b) addressed prefill-decoding interference by proposing chunked prefill, a technique that breaks the prefill stage into smaller stages to enhance batching efficiency; DistServe (Zhong et al., 2024) tackled the same issue by disaggregating the prefill and decoding stages, processing each on separate device clusters optimized for their specific stages. Splitwise (Patel et al., 2024) also disaggregated the two stages, allocating requests to distinct device clusters to achieve higher throughput and cost-effectiveness. While these systems allow users to manually specify parallel execution plans (e.g., degrees of tensor and pipeline parallelism), they do not provide guidance on determining the optimal parallel execution plan. APEX complements these works. Users can utilize APEX to simulate and derive an optimal parallel execution plan tailored to their LLM serving setup, which can then be used to configure the serving system for improved performance.
Partition-Strategy-Search Tools: Several tools have been proposed to identify an optimal configuration for LLM training. Calculon (Isaev et al., 2023) proposes a performance model that helps developers determine an optimal parallel execution plan for LLM training. vTrain (Bang et al., 2024) and ASTRA-sim (Rashidi et al., 2020) are simulators that guide users to find a system configuration that optimizes the LLM training time or the training cost. Nevertheless, these works cannot be applied to LLM serving due to the dynamism of iteration-level batching. LLMServingSim (Cho et al., 2024) and Vidur (Agrawal et al., 2024a) are the few simulation frameworks designed for LLM serving systems. Both perform fine-grained simulations that account for the dynamism of iteration-level batching. However, LLMServingSim primarily focuses on NPUs and Processing-in-Memory (PIM) architectures, as it relies on their corresponding hardware simulators to estimate execution times. In contrast, APEX utilizes operation-level profiling data to estimate execution time and does not rely on other hardware simulators. Vidur requires a model onboarding step before simulation, which involves parsing the operations within the target LLM and profiling them. APEX bypasses this requirement by capturing the key operations of LLMs and utilizing pre-collected profiling results, allowing simulations to start immediately with amortized profiling costs. More importantly, the capabilities of both works are insufficient for simulating state-of-the-art LLM serving systems. Specifically, they do not support quantization techniques, their parallelism support is limited to pipeline and tensor parallelism, and they are constrained to traditional dense LLMs, lacking provisions for emerging architectures such as Mixture of Experts (MoE) models. While it is theoretically possible to extend these simulators to accommodate additional models and parallelism strategies, the required effort remains uncertain due to the lack of evaluations on their extensibility.
@@ 6. Conlusion
In this work, we developed APEX, an extensible and dynamism-aware simulator for LLM serving. We evaluated APEX across various LLMs and distinct workloads, demonstrating its high-fidelity simulation with less than 10% relative error on average. APEX successfully identified optimal parallel execution plans, achieving up to 4.42$\times$ speedup compared to heuristic plans. The simulation proved highly efficient, providing a 71$\times$ speedup and 1234$\times$ greater cost-effectiveness than deployment on actual hardware. APEX also showcased high scalability, maintaining a similar simulation overhead when scaling from billion-scale to trillion-scale models. Additionally, we demonstrated APEX’s ease of extensibility to accommodate new models, device clusters, and more. In the future, we plan to extend APEX to support multimodal LLMs, which will involve incorporating parallel execution plans for encoders handling modalities such as vision and audio.

### v2

@@ 4.6. Beyond Identifying Optimal Execution Plan
The high simulation fidelity makes APEX applicable to other use cases beyond identifing optimal execution plans. For example, APEX can help service providers meet various service-level objectives (SLOs), such as maintaining a target time per output token (TPOT). A common strategy to meet latency requirements is to adjust the batch size by setting a maximum batch size constraint. APEX begins by simulating a subset of requests to determine an upper bound for the batch size, denoted as $m$. It then divides this upper bound into $n$ segments and simulates the request traces with various batch size constraints, ranging from $\frac{1\times m}{n}$ to $\frac{n\times m}{n}$, where $n$ is determined heuristically. Figure 9 illustrates two examples using the Llama-3.1-70B and Mistral-Large models on the Creation trace. Suppose a service provider aims to reduce TPOT to meet an SLO. Based on APEX’s results, they can estimate the required adjustment to the batch size constraint. For instance, reducing the maximum batch size from 16 to 8 yields a TPOT improvement of 18% for Llama and 14% for Mistral. However, overly restricting the maximum batch size may negatively affect performance; APEX also captures this trade-off, as shown in Figure 9, where setting the batch size constraint to 4 leads to degraded end-to-end serving latency.
5. Related Work
LLM Serving Systems: With the emergence of LLMs, numerous LLM serving systems have been proposed (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024b; Zhong et al., 2024; NVIDIA, 2024; Patel et al., 2024). Each system introduces innovations to address key challenges in LLM serving. Orca (Yu et al., 2022) introduced iteration-level batching, significantly improving serving throughput. This technique has since become a standard in LLM serving systems. vLLM (Kwon et al., 2023) proposed PagedAttention, an efficient method to manage KV cache memory usage, enabling more concurrent request batching. DistServe and Splitwise (Zhong et al., 2024; Patel et al., 2024) disaggregate the prefill and decode stages, processing stage separately (e.g., using different clusters, applying different optimizations) for better serving performance. While these systems allow users to manually specify the degree of parallelisms, they do not provide guidance on determining these configurations. APEX complements these works by automatically identifying optimal plans, which can be used to configure these serving systems.
@@ 5. Related Work
Configuration Searching Tools: LLM-Pilot (Lazuka et al., 2024) proposes a predictive model to help users select the most cost-effective hardware for serving LLMs. DynamoLLM (Stojkovic et al., 2025) proposes a framework to automatically reconfigure the cluster to optimize the energy efficiency of LLM serving. However, both LLM-Pilot and DynamoLLM only consider tensor parallelism and can be further enhanced by incorporating hybrid parallelisms, as supported by APEX. Calculon (Isaev et al., 2023) introduces a performance model to identify optimal execution plans for LLM training, while vTrain (Bang et al., 2024) and ASTRA-sim (Rashidi et al., 2020) adopt simulator-based approaches to achieve the same goal. Nevertheless, these solutions are tailored to training workloads and do not address the unique challenges of LLM serving, particularly the dynamism introduced by iteration-level batching. LLMServingSim (Cho et al., 2024) and Vidur (Agrawal et al., 2024a) are the few frameworks designed for LLM serving. Yet, LLMServingSim primarily focuses on NPUs and Processing-in-Memory architectures, and it relies on their corresponding hardware simulators to estimate execution times. In contrast, APEX utilizes operation-level profiling data to estimate execution time and does not rely on other hardware simulators. Vidur requires a model onboarding step before simulation, which involves parsing the operations within the target LLM and profiling them. APEX bypasses this requirement by capturing the key operations of LLMs and utilizing pre-collected profiling results, allowing simulations to start immediately with amortized profiling costs. In addition, neither LLMServingSim nor Vidur supports energy-aware optimization, a critical requirement for LLM service providers. They also lack support for emerging LLM serving needs, such as Mixture-of-Experts models, quantization formats, and multi-node clusters.
@@ 6. Conlusion
In this work, we developed APEX, an extensible and dynamism-aware simulator for LLM serving. We evaluated APEX across various LLMs and distinct workloads, demonstrating its high-fidelity simulation with only a 10.7% relative error on the average. APEX successfully identified latency-optimal execution plans that achieve up to 3.37$\times$ speedup compared to heuristic plans, and also energy-optimal plans that reduce energy consumption by up to 45% compared to latency-optimal plans. The simulation proved highly efficient, providing a 71$\times$ speedup and 1234$\times$ greater cost-effectiveness than deployment on actual hardware. APEX maintained a similar simulation overhead when scaling from billion-scale to trillion-scale models. While APEX already supports a broad range of LLMs, we further demonstrated its ease of extensibility to accommodate new models, device clusters, batching algorithms, and parallelisms. In the future, we plan to extend APEX to support multimodal LLMs, which will involve incorporating parallel execution plans for encoders handling modalities such as vision and audio.